# Targeted Pathway Inference for Biological Knowledge Bases via Graph Learning and Explanation

### v2

@@ Introduction
Many researchers have formulated this bio-network inference as a graph learning problem. In this setting, interactions in a bio-network are modeled as graph edges, and experimental data are embedded as node features. Various computational and machine learning methods have been proposed to infer meaningful targeted graph structures. Computational methods often rely on statistical node-centric metrics (Nacher and Akutsu 2016) to evaluate the importance of nodes. Edges connected to highly ranked nodes are considered more important. However, such objectives lack explicit inference of interactions and are computationally intractable for large bio-networks (Yang et al. 2024). Machine learning methods, particularly graph neural networks (GNNs), define explicit objectives such as link prediction or graph reconstruction, enabling direct inference of network structure (Ravindra et al. 2020). Importantly, experimental data influence the learning process via node feature aggregation, making the inferred interactions more specific to the dataset. However, existing methods are still in the early stages of exploration, and several key limitations remain unaddressed.
@@
- Topology-driven Methods. They use statistical metrics on structural properties of graphs, such as node degrees (Karger 1994; Kashtan et al. 2004a), centrality (Haynes, Hedetniemi, and Slater 2013; Wang et al. 2014; Nacher and Akutsu 2016), betweenness, or PageRank scores (Iván and Grolmusz 2011; Maehara et al. 2014) to infer which substructures exhibit a more significant influence on the overall topology, thereby identifying more targeted interactions.
- Deep Graph Learning Methods. They incorporate experimental data during the learning process by embedding data as node representations. They train GNNs with suitable objectives, such as link prediction or graph reconstruction (Yue et al. 2020; Zhang et al. 2021; Muzio, O’Bray, and Borgwardt 2021), and the links that contribute most to these objectives can be considered the targeted interactions. For instance, the works of (Hamilton, Ying, and Leskovec 2017; Gligorijević et al. 2021; Cheng et al. 2021) have been validated to predict protein functions within protein-protein interaction (PPI) networks. Moreover, GNN models have been applied to incorporate RNA-Seq data for tasks like predicting disease states and cell-cell relationships (Wang et al. 2021; Ravindra et al. 2020).
@@
Problem 2 (Pathway modeling). Many biological functions arise from long, multi-step reaction pathways that span several hops in $\mathcal{G}$. Capturing long-range dependencies is essential: (1) for functional prediction, as perturbation effects often propagate across distant nodes; and (2) for mechanistic insight into causal pathways beyond local interactions. Different experimental conditions on the same network, therefore yield distinct, data-specific subgraphs $\hat{\mathcal{G}}$, each revealing the targeted pathway most responsible for the given data. Hence,
@@
PathMamba integrates the Graph Isomorphism Network (GIN) with a novel pathway-wise Mamba model. It leverages the strengths of both global selective modeling mechanisms and message-passing GNNs. Specifically, inspired by GPS (Rampášek et al. 2022), our model avoids early-stage information loss that could arise from using GNNs in the initial layers. We employ novel pathway-wise global aggregation in efficient combination with random pathway sampling and sequential Mamba (Gu and Dao 2024) modeling.
Positional Encoding.
At each layer, node and edge features are updated by aggregating the outputs of a pathway-wise Mamba as:
@@
where $N$ is the number of nodes in the graph, and $\bf{q}^{i}$ represents the sampled pathway for node $v_{i}$. Each pathway $\bf{q}^{i}$ is a sequence of nodes $\{v_{i},v_{i_{1}},v_{i_{2}},\dots,v_{i_{L}}\}$, sampled according to a random walk process (Tonshoff et al. 2023). The sampling process $\text{Pathway}(v_{i},L)$ involves selecting a sequence of connected nodes starting from $v_{i}$. The selection of each subsequent node is determined probabilistically, guided by the graph’s adjacency structure.
@@
where $\|M\|$ encourages sparsity in the edge selection, $Y$ is a random variable representing labels {1, 2, …, C}, and $\lambda$ balances the trade-off between the classification loss and the sparsity regularization. Hence, the identified important subgraphs and node features that contribute most to specific bio-networks are considered as targeted pathways.
@@
lemmaexprforexpl(Expressiveness for explanations). When combined with higher expressive models (e.g., it distinguishes more graphs), PathExplainer can generate more finely differentiated (and potentially more “faithful”) explanation pathways (subgraphs). In contrast, less expressive models merge different graphs into larger equivalence classes, leading to non-unique, less granular explanations.
In Appendix B, we prove this by showing that the expressiveness of the underlying graph classifier $f$ determines the granularity of equivalence classes, with more expressive models enabling finer distinctions between graphs.
@@
The proof of this theorem directly comes from the recent work (Tonshoff et al. 2023; Behrouz and Hashemi 2024). They prove a similar theorem using 1-d CNN (Tonshoff et al. 2023) or SSM (Behrouz and Hashemi 2024) with randomly sampled subgraph. ∎
@@
Baselines and Metrics. We collected baselines from both message-passing GNNs and more advanced graph models, including GCN (Kipf and Welling 2017), GraphSAGE (Hamilton, Ying, and Leskovec 2017), GAT (Veličković et al. 2018), GIN (Xu et al. 2019), GPS (Rampášek et al. 2022), and Graph-Mamba (Wang et al. 2024). The detailed introduction of these baselines and the selection motivation can be found in Appendix E. We employed precision, recall, and overall accuracy for the performance evaluation. We used 650M ESM-2 for PathMamba and all baselines as the node feature encoding model.
Results. Table 1 demonstrates that PathMamba achieves the highest accuracy (0.754), outperforming all GNNs, GPS (0.726), GraphMamba (0.723). Furthermore, it secures best or second-best positions across all functional categories, demonstrating its robust ability to generalize across diverse pathway structures. The gray-shaded rows indicate the results of removing ESM-2 and modifying the model size in terms of F1 scores. When ESM-2 is removed, the accuracy decreases significantly (0.75 → 0.44). The results highlight the importance of AA-seq and the limitations of prior studies that were unable to leverage this information.
@@
Baselines. We collected baselines from three categories (i) conventional statistical methods: Random Sampling (Kashtan et al. 2004b), PersonalizedPageRank (Iván and Grolmusz 2011), and MinimumDominatingSet (Nacher and Akutsu 2016; Wuchty 2014)); (ii) gradient-based methods: Saliency (Simonyan, Vedaldi, and Zisserman 2013), InputXGradient (Shrikumar, Greenside, and Kundaje 2017), Deconvolution (Mahendran and Vedaldi 2016), ShapleyValueSampling (Strumbelj and Kononenko 2010), and GuidedBackpropagation (Springenberg et al. 2014); (iii) GNN-specific explainer method: GNNExplainer (Ying et al. 2019) and PGExplainer (Luo et al. 2020). All details can be found in Appendix E.
@@
Proposed Evaluation Metrics. We designed experiments centered on Gene Ontology (GO) analysis (Ashburner et al. 2000), focusing on the nodes within the extracted subgraphs. The results provide a list of GO terms highlighting the biological functions most significantly represented in the input gene (corresponding to protein) nodes (Ashburner et al. 2000). Then we proposed Number of Enriched Biological Functions (#EBF) and Enrichment Contribution Score (ECS) to evaluate breadth and depth of the extracted functions (Yang et al. 2024). A higher #EBF indicates broader functional diversity within the subgraph. ECS evaluates the relative contribution of the top-weighted genes. The details and definitions of metrics can be found in Appendix F.
@@
Figure 4: UpSet plot of enriched GO terms across four pathway classes, based on top feature sets from subgraphs for different methods. Orange indicates GO terms uniquely enriched in one class, and blue represents GO terms shared across multiple classes. RSS, MDS, and PPR stand for Random Sampling, Minimum Dominating Set, and Personalized PageRank, respectively.
Figure 5: Comparison of subgraphs extracted from the TCR signaling pathway using two different methods. TCR Subgraph A is from the RSS method, and TCR Subgraph B is from the proposed method. The subgraph nodes and their signaling modules are colored in red. The disruptions within signaling paths are marked in green boxes. A full pathway graph can be found in Appendix G.
@@ Appendix B Proof
$\mathcal{L}\bigl{(}\hat{\mathcal{G}};f,\mathcal{G}\bigr{)}\;=\;\alpha\,D\bigl{(}f(\mathcal{G}),f(\hat{\mathcal{G}})\bigr{)}\;+\;\beta\,\Omega\bigl{(}\hat{\mathcal{G}}\bigr{)}$,
@@ Appendix E Baselines
  • Message-passing GNNs: Graph Convolutional Network (GCN) (Kipf and Welling 2017) serves as a foundational GNN, leveraging spectral graph theory for node feature aggregation. GraphSAGE (Hamilton, Ying, and Leskovec 2017) improves scalability by introducing neighborhood sampling and learnable aggregation functions. Graph Attention Network (GAT) (Veličković et al. 2018) incorporates attention mechanisms to assign different importance to neighbors during feature aggregation. Graph Isomorphism Network (GIN) (Xu et al. 2019) achieves high expressivity, distinguishing graph structures with a focus on injective neighborhood aggregation.
@@
  • Statistical methods: Random Sampling (RRS) (Kashtan et al. 2004b) serves as a simple baseline by selecting nodes or edges randomly for comparison. Personalized PageRank (PPR) (Iván and Grolmusz 2011) computes node importance by incorporating a teleportation mechanism that biases the random walk towards specific nodes, effectively capturing both local and global graph structures. Minimum Dominating Set (MDS) (Nacher and Akutsu 2016; Wuchty 2014) identifies a minimal set of nodes that can collectively influence or dominate all other nodes in the graph, providing insights into critical nodes for coverage or control. Notably these three statistical methods only do not use AA sequence node features.
  • Gradient-based methods: Saliency (Simonyan, Vedaldi, and Zisserman 2013) highlights features based on the magnitude of input gradients. InputXGradient (Shrikumar, Greenside, and Kundaje 2017) combines input features with their gradients to capture feature significance. Deconvolution (Mahendran and Vedaldi 2016) focuses on reconstructing important input features, emphasizing positive influences. ShapleyValueSampling (Strumbelj and Kononenko 2010) estimates feature importance using a game-theoretic approach. GuidedBackpropagation (Springenberg et al. 2014) refines gradients to highlight only relevant activations.
@@ Appendix F Metrics
As a classic method of biological functional enrichment analysis, GO analysis evaluates whether specific biological processes, molecular functions, or cellular components are statistically overrepresented in a given set of genes (i.e., gene nodes from subgraphs) compared to a background gene set (Huang, Sherman, and Lempicki 2009). The results provide a list of GO terms that highlight the biological functions most significantly represented in the input gene nodes (Ashburner et al. 2000). In our study, we conducted GO enrichment analysis using the R package clusterProfiler (Yu et al. 2012) to identify enriched GO terms.
@@ H.1 T cell receptor (TCR) signaling pathway
Precise modulation of these signaling cascades is critical for maintaining immune homeostasis, as dysregulation is implicated in a spectrum of immune disorders, including autoimmunity, primary immunodeficiencies, and T cell malignancies, where aberrant activation or attenuation of TCR signaling disrupts immune tolerance, facilitates chronic inflammation, or drives oncogenic transformation. Analysis of the molecular intricacies of TCR signaling helps therapeutic interventions, including immune checkpoint modulation, CAR-T (Rafiq, Hackett, and Brentjens 2020) cell engineering, and small-molecule inhibitors aimed at restoring immune balance and targeting immune-related diseases.

### v3

@@ Introduction
Many researchers have formulated this bio-network inference as a graph learning problem. In this setting, interactions in a bio-network are modeled as graph edges, and experimental data are embedded as node features. Various computational and machine learning methods have been proposed to infer meaningful targeted graph structures. Computational methods often rely on statistical node-centric metrics (Nacher and Akutsu 2016) to evaluate the importance of nodes. Edges connected to highly ranked nodes are considered more important. However, such objectives lack explicit inference of interactions and are computationally intractable for large bio-networks (Yang et al. 2025). Machine learning methods, particularly graph neural networks (GNNs), define explicit objectives such as link prediction or graph reconstruction, enabling direct inference of network structure (Ravindra et al. 2020). Importantly, experimental data influence the learning process via node feature aggregation, making the inferred interactions more specific to the dataset. However, existing methods are still in the early stages of exploration, and several key limitations remain unaddressed.
@@
- Topology-driven Methods. They use statistical metrics on structural properties of graphs, such as node degrees (Karger 1994; Kashtan et al. 2004b), centrality (Haynes et al. 2013; Wang et al. 2014; Nacher and Akutsu 2016), betweenness, or PageRank scores (Iván and Grolmusz 2011; Maehara et al. 2014) to infer which substructures exhibit a more significant influence on the overall topology, thereby identifying more targeted interactions.
- Deep Graph Learning Methods. They incorporate experimental data during the learning process by embedding data as node representations. They train GNNs with suitable objectives, such as link prediction or graph reconstruction (Yue et al. 2020; Zhang et al. 2021; Muzio et al. 2021), and the links that contribute most to these objectives can be considered the targeted interactions. For instance, the works of (Hamilton et al. 2017; Gligorijević et al. 2021; Cheng et al. 2021) have been validated to predict protein functions within protein-protein interaction (PPI) networks. Moreover, GNN models have been applied to incorporate RNA-Seq data, for tasks like predicting disease states and cell-cell relationships (Wang et al. 2021; Ravindra et al. 2020).
@@
Problem 2 (Pathway modeling). Many biological functions arise from long, multi-step reaction pathways that span several hops in $\mathcal{G}$. Capturing long-range dependencies is essential: (1) for functional prediction, as perturbation effects often propagate across distant nodes; and (2) for mechanistic insight into causal pathways beyond local interactions. Different experimental conditions on the same network, therefore yield distinct, data-specific subgraphs $\hat{\mathcal{G}}$, each revealing targeted pathway most responsible for the given data. Hence,
@@
PathMamba integrates the Graph Isomorphism Network (GIN) with a novel pathway-wise Mamba model. It leverages the strengths of both global selective modeling mechanisms and message-passing GNNs. Specifically, inspired by GPS (Rampášek et al. 2022), our model avoids early-stage information loss that could arise from using GNNs in the initial layers. We employ novel pathway-wise global aggregation in efficient combination with random pathway sampling and sequential Mamba (Gu and Dao 2024) modeling. At each layer, node and edge features are updated by aggregating the outputs of a pathway-wise Mamba as:
@@
where $N$ is the number of nodes in the graph, and $\bf{q}^{i}$ represents the sampled pathway for node $v_{i}$. Each pathway $\bf{q}^{i}$ is a sequence of nodes $\{v_{i},v_{i_{1}},v_{i_{2}},\dots,v_{i_{L}}\}$, sampled according to a random walk process (Tonshoff et al. 2023). The sampling process $\text{Pathway}(v_{i},L)$ involves selecting a sequence of connected nodes starting from $v_{i}$. The selection of each subsequent node is determined probabilistically, guided by the graph adjacency structure.
@@
where $\|M\|$ encourages sparsity in the edge selection, $Y$ is a random variable representing labels {1, 2, …, C}, and $\lambda$ balances the trade-off between the prediction consistency and the sparsity regularization. Hence, the identified important subgraphs and node features that contribute most to specific bio-networks are considered as targeted pathways.
@@
lemmaexprforexpl(Expressiveness for explanations). When combined with higher expressive models (e.g., it distinguishes more graphs), PathExplainer can generate more finely differentiated (and potentially more “faithful”) explanation pathways (subgraphs). In contrast, a less expressive models merges different graphs into larger equivalence classes, leading to non-unique, less granular explanations.
In Appendix B, we prove this by showing that the expressiveness of the underlying a graph classifier $f$ determines the granularity of equivalence classes, with more expressive models enabling finer distinctions between graphs.
@@
The proof of this theorem directly comes from the recent work (Tonshoff et al. 2023; Behrouz and Hashemi 2024). They prove a similar theorem using 1-d CNNs (Tonshoff et al. 2023) or SSM (Behrouz and Hashemi 2024) with randomly sampled subgraphs. Since our method adopts Mamba (an SSM architecture) combined with the random sampling strategy, their theoretical results are directly applicable to our setting. ∎
@@
Baselines and Metrics. We collected baselines from both message-passing GNNs and more advanced graph models, including GCN (Kipf and Welling 2017), GraphSAGE (Hamilton et al. 2017), GAT (Veličković et al. 2018), GIN (Xu et al. 2019), GPS (Rampášek et al. 2022), and Graph-Mamba (Wang et al. 2024). The detailed introduction of these baselines and the selection motivation can be found in Appendix E. We employed precision, recall, and overall accuracy for the performance evaluation. We used 650M ESM-2 for PathMamba and all baselines as the node feature encoding model.
Results. Table 1 demonstrates that PathMamba achieves the highest accuracy (0.744), outperforming all GNNs, GPS (0.726), GraphMamba (0.723). Furthermore, it secures best or second-best positions across all functional categories, demonstrating its robust ability to generalize across diverse pathway structures. The gray-shaded rows indicate the results of removing ESM-2 and modifying the model size in terms of F1 scores. When ESM-2 is removed, the accuracy decreases significantly (0.74 → 0.44). The results highlight the importance of AA-seq and the limitations of prior studies that were unable to leverage this information.
@@
Baselines. We collected baselines from three categories (i) conventional statistical methods: Random Sampling (Kashtan et al. 2004a), PersonalizedPageRank (Iván and Grolmusz 2011), and MinimumDominatingSet (Nacher and Akutsu 2016; Wuchty 2014)); (ii) gradient-based methods: Saliency (Simonyan et al. 2013), InputXGradient (Shrikumar et al. 2017), Deconvolution (Mahendran and Vedaldi 2016), ShapleyValueSampling (Strumbelj and Kononenko 2010), and GuidedBackpropagation (Springenberg et al. 2014); (iii) GNN-specific explainer method: GNNExplainer (Ying et al. 2019) and PGExplainer (Luo et al. 2020). All details can be found in Appendix E.
@@
Proposed Evaluation Metrics. We designed experiments centered on Gene Ontology (GO) analysis (Ashburner et al. 2000), focusing on the nodes within the extracted subgraphs. The results provide a list of GO terms highlighting the biological functions most significantly represented in the input gene (corresponding to protein) nodes (Ashburner et al. 2000). Then we proposed Number of Enriched Biological Functions (#EBF) and Enrichment Contribution Score (ECS) to evaluate breadth and depth of the extracted functions (Yang et al. 2025). A higher #EBF indicates broader functional diversity within the subgraph. ECS evaluates the relative contribution of the top-weighted genes. The details and definitions of metrics can be found in Appendix F.
@@
Figure 4: UpSet plot of enriched GO terms across four pathway classes, based on top feature sets from subgraphs for different methods. Orange indicates GO terms uniquely enriched in one class, and blue represents GO terms shared across multiple classes. MDS, and PPR stand for Minimum Dominating Set, and Personalized PageRank, respectively.
Figure 5: Comparison of subgraphs extracted from the TCR signaling pathway using two different methods. TCR Subgraph A is from the RSS method, and TCR Subgraph B is from the proposed method. The subgraph nodes and their signaling modules are colored in red. The disruptions within signaling paths are marked in green boxes. A full pathway graph can be found in Appendix H.
@@ Conclusion
Acknowledgments
The authors would like to thank the reviewers. This work was supported by JST BOOST, Japan Grant Number JPMJBS2402, “Program for Leading Graduate Schools” of The University of Osaka, JSPS KAKENHI Grant-in-Aid for Scientific Research Number JP24K20778, NSF award SCH-2205289, SCH-2014438, and IIS-2034479, JST CREST JPMJCR23M3, JST START JPMJST2553, JST CREST JPMJCR20C6, JST K Program JPMJKP25Y6, JST COI-NEXT JPMJPF2009, JST COI-NEXT JPMJPF2115, the Future Social Value Co-Creation Project - The University of Osaka.
@@ Appendix B Proof
[…]
@@ Appendix E Baselines
  • Message-passing GNNs: Graph Convolutional Network (GCN) (Kipf and Welling 2017) serves as a foundational GNN, leveraging spectral graph theory for node feature aggregation. GraphSAGE (Hamilton et al. 2017) improves scalability by introducing neighborhood sampling and learnable aggregation functions. Graph Attention Network (GAT) (Veličković et al. 2018) incorporates attention mechanisms to assign different importance to neighbors during feature aggregation. Graph Isomorphism Network (GIN) (Xu et al. 2019) achieves high expressivity, distinguishing graph structures with a focus on injective neighborhood aggregation.
@@
  • Statistical methods: Random Sampling (RRS) (Kashtan et al. 2004a) serves as a simple baseline by selecting nodes or edges randomly for comparison. Personalized PageRank (PPR) (Iván and Grolmusz 2011) computes node importance by incorporating a teleportation mechanism that biases the random walk towards specific nodes, effectively capturing both local and global graph structures. Minimum Dominating Set (MDS) (Nacher and Akutsu 2016; Wuchty 2014) identifies a minimal set of nodes that can collectively influence or dominate all other nodes in the graph, providing insights into critical nodes for coverage or control. Notably these three statistical methods only do not use AA sequence node features.
  • Gradient-based methods: Saliency (Simonyan et al. 2013) highlights features based on the magnitude of input gradients. InputXGradient (Shrikumar et al. 2017) combines input features with their gradients to capture feature significance. Deconvolution (Mahendran and Vedaldi 2016) focuses on reconstructing important input features, emphasizing positive influences. ShapleyValueSampling (Strumbelj and Kononenko 2010) estimates feature importance using a game-theoretic approach. GuidedBackpropagation (Springenberg et al. 2014) refines gradients to highlight only relevant activations.
@@ Appendix F Metrics
As a classic method of biological functional enrichment analysis, GO analysis evaluates whether specific biological processes, molecular functions, or cellular components are statistically overrepresented in a given set of genes (i.e., gene nodes from subgraphs) compared to a background gene set (Huang et al. 2009). The results provide a list of GO terms that highlight the biological functions most significantly represented in the input gene nodes (Ashburner et al. 2000). In our study, we conducted GO enrichment analysis using the R package clusterProfiler (Yu et al. 2012) to identify enriched GO terms.
@@ H.1 T cell receptor (TCR) signaling pathway
Precise modulation of these signaling cascades is critical for maintaining immune homeostasis, as dysregulation is implicated in a spectrum of immune disorders, including autoimmunity, primary immunodeficiencies, and T cell malignancies, where aberrant activation or attenuation of TCR signaling disrupts immune tolerance, facilitates chronic inflammation, or drives oncogenic transformation. Analysis of the molecular intricacies of TCR signaling helps therapeutic interventions, including immune checkpoint modulation, CAR-T (Rafiq et al. 2020) cell engineering, and small-molecule inhibitors aimed at restoring immune balance and targeting immune-related diseases.